If explict costs are $160,000 and implict costs are $72,000, economic profit is accounting profit by
Answer #1

Explicit cost = $160,000

Implicit cost = $72,000

Assume Revenue = $X

Accounting Profit = Revenue - Explicit cost = X - 160,000

Economic Profit = Revenue - Explicit cost - Implicit Cost = X - 160,000 - 72,000 = X - 232,000